# PreGIP: Watermarking the Pretraining of Graph Neural Networks for Deep Intellectual Property Protection

### v1

@@ Appendix C Proofs
Appendix D Additional Details of Experimental Settings
D.1 Datasets and Learning Settings
Table 3: Dataset statistics for semi-supervised learning
Datasets
#Graphs
#Avg. Nodes
#Avg. Edges
#Features
#Classes
#Tasks
PROTEINS
1,113
39.1
145.6
3
2
1
NCI1
4,110
29.9
32.3
38
2
1
FRANKENSTEIN
4337
16.9
17.9
780
2
1
Semi-Supervised Setting. We first conduct experiments on widely used benchmark datasets in the setting of semi-supervised (You et al., 2020) on graph classification via pre-training & fine-tuning where a small set of graphs are provided with labels for fine-tuning. Specifically, three public benchmarks, i.e., PROTEINS, NCI1 (Sergey et al., 2020) and FRANKENSTEIN (Orsini et al., 2015), are used in the semi-supervised setting. PROTEINS is a set of proteins that are classified as enzymes or non-enzymes. NCI1 is a molecular graph set, which is relative to anti-cancer screens where the chemicals are assessed as positive or negative to cell lung cancer. FRANKENSTEIN is molecule dataset for mutagenicity (AMES) classification. The statistics details of these datasets are summarized in Tab. 3. Additionally, for pre-training, the epoch number is set as $1000$ and the learning rate is set as $0.0003$. During fine-tuning, the label rate is set as $50\%$. We conduct experiments across various train/test splits in the downstream tasks, repeating them 50 times to derive the mean and standard deviation of the results.
Table 4: Dataset statistics for transfer learning
Utilization
Datasets
#Graphs
#Avg. Nodes
#Avg. Edges
#Features
#Classes
#Tasks
Pretraining
ZINC15
200,000
23.1
24.8
120
0
0
Finetune
Tox21
7,831
18.6
38.6
9
2
12
ToxCast
8,597
18.7
38.4
9
2
617
BBBP
2,050
23.9
51.6
9
2
1
BACE
1,513
34.1
73.7
9
2
1
SIDER
1,427
33.6
70.7
9
2
27
ClinTox
1,484
26.1
55.5
9
2
2
Transfer Learning Setting. Following our semi-supervised learning experiments, we proceed to conduct experiments in a transfer learning setting (Hu et al., 2020). We utilize 200,000 unlabeled molecules sampled from the ZINC15 dataset (Sterling & Irwin, 2015) for pre-training, and involve 6 public molecular graph classification benchmarks in the downstream tasks. The statistics details of these datasets are summarized in Tab. 4. Especially, given that Tox21, ToxCast, SIDER, and ClinTox encompass multiple tasks, we compute the final IP Gap and IP ROC for these datasets by averaging the IP Gap and IP ROC across their respective tasks. During pre-training on ZINC15, the epoch number is set as $200$ and the learning rate is $0.001$. The label rate in fine-tuning is set as $50\%$. Similar to the learning setting of semi-supervised learning, we conduct experiments across various train/test splits in the downstream tasks for $10$ times and report the average and standard deviation of the results.
D.2 Details of Baselines
In this section, we present the details of four state-of-the-art watermarking methods as baselines to protect the IP of the pretraining of GNNs:
  • Deepsign (Rouhani et al., 2018): This is a watermarking-based method for protecting the ownership of deep learning models in the image domain. The main idea is to ensure the watermark samples to have the same embeddings. During the verification phase, this method requires to access to the intermediate representation of the piracy model. For a watermarked model, the embeddings of watermark samples will be similar, resulting to the same predictions from the downstream classifier. Hence, to extend Deepsign to the black-box verification, the fraction that watermark samples are predicted as the same class is used for IP indication score.
  • RWM-Pre (Zhao et al., 2021): This baseline is inspired by RWM (Zhao et al., 2021), which focuses on prompting watermarked models to predict Erdos-Renyi (ER) random graphs with random node features and labels to certain classes. Given that RWM is primarily designed for node classification in semi-supervised learning, in order to adapt it to our GNN pretraining scenario, we pretrain a GIN encoder to learn representations. Then, we apply a clustering algorithm to these learned representations, using the resulting cluster labels as pseudo classes for RWM. This approach is denoted as RWM-Pre. During the verification phase, RWM-Pre will compute the fraction of watermark graphs predicted as the same class, serving as the IP indication score.
  • CWM-Pre (Bansal et al., 2022): This is a certifiable watermarking method to provide watermark certification by injecting noises to model parameters during the watermarking. This method is based on (Zhang et al., 2018) which predicts special samples as certain class. Therefore, to adapt this method for watermarking GNN pretraining, we incorporate the noise injection technique from (Bansal et al., 2022) into RWM-Pre, establishing the baseline as CWM-Pre.
  • GCBA (Zhang et al., 2023): It is a backdoor attack method for graph contrastive learning. In our setting, we apply the poisoning-based GCBA that aims to align the embeddings of trigger-attached graphs to those of the target class. To evaluate the IP performance of GCBA, we follow the setting in (Adi et al., 2018) to quantify the difference between the attack success rate of trigger-attached graphs and the classification accuracy of the clean test graphs, both before and after backdooring. These metrics are referred to as the IP Gap and accuracy, respectively, as detailed in Sec. 4.1.
D.3 Implementation Details
In this paper, we follow the settings in (Hu et al., 2020), where we initially train the GNN encoders using GCL or other pretraining methods (Hu et al., 2020). Then, following (Hu et al., 2020), we employ the obtained embeddings to train and evaluate a classifier for the downstream tasks. A $2$-layer GIN is deployed as the backbone GNN encoder. The hidden dimension is set as 256 for all experiments. GraphCL is set as the target GCL methods, which is implemented based on PyGCL library (Zhu et al., 2021). The edge prediction and edge masking strategies are implemented based on the source code published by the original authors  [^1] https://github.com/snap-stanford/pretrain-gnns/. The hyperparameter settings of the pretraining are the same as the cited paper.
In the watermark graph pair generation, the probability of connecting each pair of nodes is set as 0.2. We maintain a consistent difference of 15 nodes in the node count between watermark graphs within each pair. By leveraging the randomness in synthetic structures and varying the sizes of the graphs, we are able to generate distinct synthetic graphs to build multiple pairs watermark graphs. In the watermarking loss, the $Q$ and $m$ in Eq.(2) are set as 256 and 1.0.
In experiments of PreGIP, we configure the number of watermark graph pairs as $20$ and set $\lambda=1$, $\epsilon=2$. To make fair comparisons, all the hyperparameters of these methods are tuned based on the watermarking loss and the pretraining loss on validation set. All models undergo training on an A6000 GPU with 48G memory.
Appendix E Additional Results of Watermarking Pretraining
E.1 Hyperparameter Analysis
(a) IP ROC (%)
(b) Accuracy (%)
Figure 6: Hyperparameter analysis on PROTEINS.
We further investigate how hyperparameters $\lambda$ and $\epsilon$ affect the performance of PreGIP, where $\lambda$ controls the contributions of the proposed watermarking loss on the pretrained models, and $\epsilon$ controls the magnitude of changes to the model parameters $\theta$ in finetuning-resist watermark injection. More specifically, we vary $\lambda$ and $\epsilon$ as $\{0.01,0.1,1,5,10\}$ and $\{0.1,1,2,5,10\}$, respectively. We report the IP ROC and accuracy on PROTEINS datasets. The experimental seting is semi-supervised under the finetuning scenario. All other settings are the same as Appendix D.3. Similar trends are also observed on other datasets and pretraining methods. The results are shown in Fig. 6. From this figure, we find that IP ROC will increase with the increasing of $\lambda$ and $\epsilon$. As for the accuracy of the watermarked GNN encoder in downstream tasks, it will firstly stay consistent then decrease if the $\lambda$ and $\epsilon$ is overly high. From this figure, we can find that when the $\epsilon$ is set around $2$ and $\lambda$ ranges from 0.1 to 1, we can achieve the balance between the GNN encoder utility and watermarking performance. Specifically, according to the hyperparameter analysis on PROTEINS, we set $\epsilon$ and $\lambda$ as 2 and 1. Then, the same hyperparameter setting is applied for all other datasets and experimental settings without further tuning.
E.2 Additional Results of Flexibility of Watermarking
The full comparison results are reported in Tab. 5. The observations are rather similar to those of Fig. 3 in Sec. 4.2.
Table 5: Results in protecting IP of edge prediction-pretrained GNNs (Hu et al., 2020) in the semi-supervised setting.
Scenario
Dataset
Metrics (%)
Non-Watermarked
Deepsign
RWM-Pre
CWM-Pre
PreGIP
Fix PROTEINS Accuracy
68.0$\pm$1.1
66.6$\pm$1.4
62.3$\pm$1.2
62.9$\pm$0.9
69.1$\pm$1.1
IP Gap
6.0$\pm$4.7
29.5$\pm$9.7
8.1$\pm$2.5
31.0$\pm$3.5
IP ROC
34.5
90.0
58.0
100
NCI1 Accuracy
68.7$\pm$0.9
69.0$\pm$1.0
68.4$\pm$0.8
68.5$\pm$0.8
69.7$\pm$0.6
IP Gap
7.5$\pm$0.1
12.5$\pm$1.3
11.0$\pm$2.3
43.0$\pm$9.8
IP ROC
84.0
78.0
82.5
100
Finetune PROTEINS Accuracy
68.2$\pm$1.5
68.2$\pm$2.2
67.5$\pm$1.6
63.1$\pm$3.1
67.9$\pm$2.2
IP Gap
3.0$\pm$1.0
6.0$\pm$5.0
14.0$\pm$7.5
35.0$\pm$1.4
IP ROC
54.0
61.0
58.5
87.5
NCI1 Accuracy
69.6$\pm$0.8
68.8$\pm$0.8
69.7$\pm$0.9
69.7$\pm$0.6
70.8$\pm$0.9
IP Gap
4.5$\pm$0.7
5.0$\pm$1.5
9.5$\pm$3.4
16.5$\pm$14.7
IP ROC
43.0
60.5
68.0
85.0
E.3 Additional Results of Pretraining for Transfer Learning
The comparison results of the edge masking pretraining on 6 downstream datasets are reported in Tab. 6. The observations are rather similar to those of Tab. 2 in Sec. 4.2.
Table 6: Comparison with baselines in protecting IP of edge masking-pretrained GNNs in the transfer learning setting.
Dataset
Metrics (%)
Non-Watermarked
Deepsign
RWM-Pre
CWM-Pre
PreGIP
Tox21
Accuracy
92.7$\pm$0.2
92.8$\pm$0.1
92.9$\pm$0.2
92.8
92.9$\pm$0.2
IP Gap
4.9$\pm$11.6
7.5$\pm$4.8
11.9$\pm$10.6
32.4$\pm$28.0
IP ROC
38.0
62.0
66.0
96.0
ToxCast
Accuracy
83.5$\pm$0.3
83.7$\pm$3.2
83.4$\pm$0.4
83.3$\pm$0.1
83.8$\pm$0.1
IP Gap
4.4$\pm$3.2
27.6$\pm$24.6
7.2$\pm$2.4
33.6$\pm$12.1
IP ROC
96.0
72.0
70.0
100.0
BBBP
Accuracy
83.9$\pm$1.8
85.5$\pm$1.1
85.3$\pm$1.2
85.2$\pm$1.1
85.4$\pm$1.4
IP Gap
12.0$\pm$6.8
25.4$\pm$22.0
26.1$\pm$12.0
39.0$\pm$30.5
IP ROC
90.0
80.0
84.0
94.0
BACE
Accuracy
70.2$\pm$0.8
73.4$\pm$2.2
74.3$\pm$1.8
74.4$\pm$1.6
74.5$\pm$2.4
IP Gap
6.7$\pm$6.4
26.7$\pm$25.9
12.2$\pm$10.4
38.0$\pm$26.4
IP ROC
68.0
80.0
78.0
84.0
SIDER
Accuracy
75.3$\pm$0.3
75.1$\pm$0.3
74.8$\pm$0.4
74.7$\pm$0.3
74.8$\pm$0.2
IP Gap
3.7$\pm$3.4
27.8$\pm$27.1
21.4$\pm$17.5
31.2$\pm$29.6
IP ROC
20.0
80.0
64.0
80.0
ClinTox
Accuracy
91.3$\pm$0.5
91.6$\pm$1.2
91.6$\pm$1.7
91.4$\pm$1.3
91.8$\pm$1.2
IP Gap
2.0$\pm$1.0
37.2$\pm$34.0
33.2$\pm$31.5
45.0$\pm$42.0
IP ROC
40.0
82.0
80.0
86.0
E.4 Additional Results of Impact of the Number of Watermark Graph Pairs
The IP ROC results of PreGIP on PROTEINS and FRANKENSTEIN are reported in Tab. 7. The observations are rather similar to those of Fig. 4 in Sec. 4.3.
Table 7: IP ROC of PreGIP under different numbers of watermark graph pairs (%).
Dataset
1
5
10
20
50
100
PROTEINS
55
92.5
96.5
100
100
100.0
FRANKENSTEIN
50
90
100
100
100
100.0

### v2

@@ Appendix A Proofs
Table 4. Dataset statistics for semi-supervised learning
Dataset #Graphs #Avg. Nodes #Avg. Edges #Classes
PROTEINS 1,113 39.1 145.6 2
NCI1 4,110 29.9 32.3 2
FRANKENSTEIN 4,337 16.9 17.9 2
Table 5. Dataset statistics for transfer learning
Utilization Dataset #Graphs #Avg. Nodes #Avg. Edges
Pretraining ZINC15 200,000 23.1 24.8
Finetune Tox21 7,831 18.6 38.6
ToxCast 8,597 18.7 38.4
BBBP 2,050 23.9 51.6
BACE 1,513 34.1 73.7
SIDER 1,427 33.6 70.7
ClinTox 1,484 26.1 55.5
Appendix B Experimental Setting Details
B.1. Datasets and Learning Settings
Semi-Supervised Setting. In the semi-supervised setting, GNNs are trained via pre-training & fine-tuning. For pre-training, the epoch number is set as $1000$ and the learning rate is set as $0.0003$. During fine-tuning, the label rate is set as $50\%$. We conduct experiments across various train/test splits in the downstream tasks, repeating them 50 times to derive the average results.
Transfer Learning Setting. In the transfer learning setting (Hu et al., 2020). We utilize 200,000 unlabeled molecules sampled from the ZINC15 dataset (Sterling and Irwin, 2015) for pre-training, and involve 6 public molecular graph classification benchmarks in the downstream tasks. The statistics details of these datasets are summarized in Tab. 5. During pre-training, the epoch number is set as $200$ and the learning rate is $0.001$. The label rate in fine-tuning is set as $50\%$. We conduct experiments across various train/test splits in the downstream tasks for $10$ times and report the average results.
Table 6. Results in watermarking edge prediction-pretrained GNNs (Hu et al., 2020) under semi-supervised setting.
Scenario Dataset Metrics (%) Non-Watermarked WPE SSL-WM Deepsign RWM-Pre CWM-Pre PreGIP
Fix PROTEINS Accuracy 68.0$\pm$1.1 66.9$\pm$1.5 66.3$\pm$1.5 66.6$\pm$1.4 62.3$\pm$1.2 62.9$\pm$0.9 69.1$\pm$1.1
IP Gap - 5.8$\pm$1.5 3.7$\pm$1.9 6.0$\pm$4.7 29.5$\pm$9.7 8.1$\pm$2.5 31.0$\pm$3.5
IP ROC - 58.6 71.2 65.5 90.0 58.0 100
NCI1 Accuracy 68.7$\pm$0.9 65.7$\pm$1.6 67.9$\pm$0.9 69.0$\pm$1.0 68.4$\pm$0.8 68.5$\pm$0.8 69.7$\pm$0.6
IP Gap - 34.8$\pm$14.6 5.0$\pm$4.9 7.5$\pm$0.1 12.5$\pm$1.3 11.0$\pm$2.3 43.0$\pm$9.8
IP ROC - 98.4 80.0 84.0 78.0 82.5 100
Finetune PROTEINS Accuracy 68.2$\pm$1.5 67.8$\pm$1.7 67.8$\pm$1.9 68.2$\pm$2.2 67.5$\pm$1.6 63.1$\pm$3.1 67.9$\pm$2.2
IP Gap - 2.3$\pm$2.2 1.5$\pm$1.3 3.0$\pm$1.0 6.0$\pm$5.0 14.0$\pm$7.5 35.0$\pm$1.4
IP ROC - 52.2 79.0 54.0 61.0 58.5 87.5
NCI1 Accuracy 69.6$\pm$0.8 67.1$\pm$1.3 68.8$\pm$1.2 68.8$\pm$0.8 69.7$\pm$0.9 69.7$\pm$0.6 70.8$\pm$0.9
IP Gap - 39.3$\pm$38.7 0.0$\pm$0.0 4.5$\pm$0.7 5.0$\pm$1.5 9.5$\pm$3.4 16.5$\pm$14.7
IP ROC - 64.0 50.0 57.0 60.5 68.0 85.0
Table 7. Results in watermarking of edge prediction-pretrained GNNs (Hu et al., 2020) in the transfer learning setting.
Dataset Metrics (%) Non-Watermarked WPE SSL-WM Deepsign RWM-Pre CWM-Pre PreGIP
Tox21 Accuracy 92.7$\pm$0.2 92.6$\pm$0.2 92.2$\pm$0.1 92.8$\pm$0.1 92.9$\pm$0.2 92.8 92.9$\pm$0.2
IP Gap - 0.9$\pm$0.6 0.3$\pm$0.2 4.9$\pm$11.6 7.5$\pm$4.8 11.9$\pm$10.6 32.4$\pm$28.0
IP ROC - 90.0 80.0 62.0 62.0 66.0 96.0
ToxCast Accuracy 83.5$\pm$0.3 83.5$\pm$0.5 82.8$\pm$0.2 83.7$\pm$3.2 83.4$\pm$0.4 83.3$\pm$0.1 83.8$\pm$0.1
IP Gap - 2.1$\pm$0.3 2.9$\pm$2.4 4.4$\pm$3.2 27.6$\pm$24.6 7.2$\pm$2.4 33.6$\pm$12.1
IP ROC - 60.0 70.0 96.0 72.0 70.0 100.0
BBBP Accuracy 83.9$\pm$1.8 82.1$\pm$1.3 79.2$\pm$1.0 85.5$\pm$1.1 85.3$\pm$1.2 85.2$\pm$1.1 85.4$\pm$1.4
IP Gap - 2.5$\pm$1.5 7.0$\pm$3.5 12.0$\pm$6.8 25.4$\pm$22.0 26.1$\pm$12.0 39.0$\pm$30.5
IP ROC - 70.0 60.0 90.0 80.0 84.0 94.0
BACE Accuracy 70.2$\pm$0.8 68.4$\pm$0.8 68.9$\pm$0.5 73.4$\pm$2.2 74.3$\pm$1.8 74.4$\pm$1.6 74.5$\pm$2.4
IP Gap - 5.3$\pm$1.5 17.5$\pm$12.0 6.7$\pm$6.4 26.7$\pm$25.9 12.2$\pm$10.4 38.0$\pm$26.4
IP ROC - 50.0 80.0 68.0 80.0 78.0 84.0
SIDER Accuracy 75.3$\pm$0.3 75.2$\pm$0.5 75.2$\pm$0.3 75.1$\pm$0.3 74.8$\pm$0.4 74.7$\pm$0.3 74.8$\pm$0.2
IP Gap - 2.1$\pm$1.6 0.8$\pm$0.6 3.7$\pm$3.4 27.8$\pm$27.1 21.4$\pm$17.5 31.2$\pm$29.6
IP ROC - 70.0 60.0 60.0 80.0 64.0 80.0
ClinTox Accuracy 91.3$\pm$0.5 92.2$\pm$1.2 91.4$\pm$1.3 91.6$\pm$1.2 91.6$\pm$1.7 91.4$\pm$1.3 91.8$\pm$1.2
IP Gap - 2.0$\pm$1.0 0.0$\pm$0.0 2.0$\pm$1.0 37.2$\pm$34.0 33.2$\pm$31.5 45.0$\pm$42.0
IP ROC - 80.0 50.0 60.0 82.0 80.0 86.0
B.2. Details of Baselines
  • WPE (Wu et al., 2022): It enforce the encoder to have dissimilar embeddings for a trigger-attached sample with the original sample. Then, the fraction of trigger-attached samples and original samples receiving the different predictions is used as the IP indication score. To extend this approach to graph-structured data, we use randomly generated graphs as triggers in the implementation.
  • SSL-WM (Lv et al., 2022): This is a black-box watermarking approach for self-supervised learning. SSL-WM enforces the encoder to give similar embeddings to the watermark samples, which will lead to abnormal entropy of the posteriors of watermark samples. During the verification phase, fraction of abnormal posteriors of watermark samples are used as IP indication score. To extend this method to graph-structured data, we adopt randomly generated graphs as the watermark samples.
  • Deepsign (Rouhani et al., 2018): Deepsign enforces watermark samples to have similar embeddings, resulting in identical downstream predictions. Originally requiring intermediate representations, it is adapted here for black-box verification by measuring the fraction of watermark samples predicted as the same class.
  • RWM-Pre (Zhao et al., 2021): This baseline, inspired by RWM (Zhao et al., 2021), prompts models to classify random Erdos-Renyi graphs with random features into specific classes. Originally designed for semi-supervised node classification, we adapt it for GNN pretraining (RWM-Pre) by clustering pretrained GIN embeddings to generate pseudo classes. IP verification uses the fraction of watermark graphs predicted into the same class.
  • CWM-Pre (Bansal et al., 2022): A certifiable watermarking method using noise injection into model parameters to predict special samples as specific classes. We adapt it for GNN pretraining by integrating its noise injection technique into RWM-Pre, naming the approach CWM-Pre.
  • GCBA (Zhang et al., 2023): It is a backdoor attack method for graph contrastive learning. We follow the setting in (Adi et al., 2018) to quantify the difference between the attack success rate of trigger-attached graphs.
Appendix C Flexibility of Watermarking
The complete results of watermarking the edge prediction pretraining GNN encoder under semi-supervised learning and transfer learning settings are reported in Tab. 7 and Tab. 7, respectively. The observations are rather similar to those of Fig. 3 and Tab. 2 in Sec. 4.2. These results further demonstrate the effectiveness of our PreGIP in watermarking different pretraining strategies for various learning settings.